Southwest Sands currently has 23,000 shares of stock outstanding. It is considering issuing $134,000 of debt at an interest rate of 7.5 percent. The break-even level of EBIT between these two capital structure options is $76,000. How many shares of stock will be repurchased if the company undergoes the recapitalization? Ignore taxes.
Answer #1

EPS =(EBIT-Interest)/No of shares

76000/23000 =(76000-134000*.075)/x

No of shares or x =19,958.55

Shares repurchased =23000-19,958.55 =3041.45 shares